# Schwarzschild Black Hole Surrounded by Perfect Fluid Dark Matter in the presence of Quintessence Matter Field

### v1

@@ Abstract
In this manuscript, we examine the Schwarzschild black hole surrounded by perfect fluid dark matter in the presence of the quintessence matter field in the context of thermodynamics, shadows, and quasinormal modes. We find that the quintessence matter field sets an upper bound value on the horizon where dark matter settles a lower limit on the event horizon but also expands the upper bound value of the horizon. We discuss how this impact alters the black hole’s thermodynamics, shadows, and quasinormal modes.
@@ 1 Introduction
The universe is a great, mysterious puzzle. Just when we believe we have the clue to solve a piece of this puzzle, we are confronted with the reality of how little we understand. At the very beginning of this century, we thought that with the observation of the Higgs boson, the Standard Model would be roughly complete, and hence, we would be able to explain important phenomena. After the discovery of the Higgs boson, we obtained reliable clues that reveal that only about $5\%$ of our universe is composed of directly observable ordinary matter, and the rest is a combination of dark matter and dark energy [1, 2]. So far we have not directly observed these dark sector components, but through indirect findings, we have been able to discuss evidence of their existence [3, 4]. For example, in the literature, the presence of dark matter was predicted by the rotation curves of spiral galaxies [5, 6, 7], dynamical stability of disk galaxies [9, 8], large-scale structure formation of the universe [10], gravitational lensing observations by galaxies and clusters of galaxies [11, 12], cosmic microwave background [13, 14], and baryon acoustic oscillations [15]. Many theories have been put forward so far to model dark matter [16, 17, 18, 19, 20, 21]. Most of these theories go beyond the Standard Model [22, 23, 24, 25, 26, 27, 28, 29], while few of them do not [30, 31, 32, 33]. Among all these models, the perfect fluid dark matter (PFDM) model, which treats dark matter as a perfect fluid, presents interesting theoretical aspects. For example, Kiselev derived exact static spherically symmetric black hole solutions out of Einstein equations with the PFDM [34]. Later, Li et al used the logarithmic term, which arises from the PFDM model of Kiselev, to discuss the asymptotic rotation curves for the dark matter in the halo-dominated region [35]. In another work, Xu et al generalized Kiselev’s solution to Kerr-like black holes employing the Newman-Janis method and examined Kerr-(A)dS black hole properties [36]. Soon after this work, PFDM impacts on deflection angle, black hole’s shadow, naked singularity, and thermal properties were investigated in [37, 38, 39, 40]. We observe that in very recent works the features of other black holes such as Reissner-Nordström, Reissner-Nordström-AdS, Schwarzschild, Schwarzschild-dS, and non-linear magnetic-charged AdS black holes were studied by assuming PFDM surrounded them [42, 41, 44, 43, 45, 46, 47, 48].
@@
Reliable astronomical observations put forward that our universe is expanding at an accelerating rate [49, 50, 51], and dark energy, the other component of the dark sector, is thought to be behind this fact. The cosmological constant, whose observational and theoretical values differ by $120$ orders of magnitude [52], cannot explain this expansion. Theoretical physicists proposed new models using dynamical scalar fields to describe dark energy, taking into account that the cosmological constant does not alter with time [53, 59, 54, 55, 56, 57, 58, 60]. Among different models, the quintessence matter model stands in the form where the equation of state is limited by the linear relationship between pressure, energy density and a state parameter [55]. Using the quintessence matter field, Kiselev solved Einstein’s equation and obtained a general solution form of the static spherically symmetric black hole [61]. Subsequently, in the presence of quintessence matter new solutions of other black holes were obtained, and the impacts of the quintessence field on their thermal quantities, shadows, and quasinormal modes were discussed widely in [62, 63, 64, 65, 66, 67, 68, 70, 69, 71, 72, 73, 75, 74, 77, 76, 80, 78, 79, 81, 82, 83, 84, 85, 87, 86, 88, 89].
@@
Although black hole solutions have been sought for more than a century, the problem of their detection remained an open debate. Until the successful observational results of the Event Horizon Telescope were published [90, 91, 92, 93, 94], some authors suggested that not the black holes but their shadows could be detected [95, 97, 96]. According to this proposal, if a black hole is located between a light source and an observer, it should be expected that some of the light rays would be deflected by the intense gravitational field of the black hole and reach the observer. However, some photons would fall into the black hole instead of this deflection and create a dark region called a shadow. The apparent shape of the black hole is assumed to correspond to the boundary of the shadow. Nowadays, we observe many papers that examine the impact the space-time geometry, dark energy, and dark matter on shadow images in the literature [98, 99, 100, 105, 106, 116, 109, 108, 111, 112, 110, 115, 123, 119, 124, 113, 107, 101, 102, 103, 117, 118, 104, 120, 121, 129, 132, 122, 125, 130, 126, 114, 127, 128, 131, 133].
On the other hand, the detection of gravitational waves has provided new clues to solve our puzzle [134, 135]. Advanced observations of gravitational waves in particular can be employed to test many alternative theories of general relativity. Black holes, which are in their final stage, are assumed to emit gravitational waves in the form of quasinormal radiation [136]. Mathematically, quasinormal modes are considered to be in complex number forms, where the real and imaginary sections are associated with the actual oscillation and the inverse damping time, respectively [137]. Since they are independent of the initial perturbation, they are frequently assumed as the fingerprints of black holes [138]. In literature, numerous studies have been conducted by using analytic and numerical methods to delve into quasinormal modes of black holes within various scenarios [139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167].
In this contribution, we intend to explore thermodynamics, shadows, and quasinormal modes features of Schwarzschild black hole surrounded by PFDM in the presence of the quintessence matter field. To achieve our goal we organize the paper as follows: In Sec. 2, we present a brief review of the mathematical construction of the considered scenario. Then, in Sec. 3, we delve into its thermodynamics. In Sec. 4, we describe the expected shadow images. After that, in Sec. 5, we derive the quasinormal modes by using two different approaches. Finally, we conclude the manuscript in Sect. 6. Here, we have to note that in the rest of the manuscript, we assume $G=c=\hbar=k_{B}=1$.
@@ 2 Black Hole Solution
We begin with the action that describes the Schwarzschild black hole surrounded by PFDM [35, 36, 42]
$\mathcal{S}_{0}=\int d^{4}x\sqrt{-g}\left[\frac{1}{16\pi}\left(R-2\Lambda\right)-4\pi\mathcal{L}_{PFDM}\right].$ (1)
Here, $g$ and $R$ denote the determinant of the metric tensor and the scalar curvature, respectively. Then, we consider the quintessence matter field to model the impact of dark energy [168]
$\mathcal{S}=\mathcal{S}_{0}+\int d^{4}x\sqrt{-g}\mathcal{L}_{q},$ (2)
@@
where the quintessence field Lagrange density is given in terms of the quintessential scalar field, $\phi$, and the potential, $V\left(\phi\right)$, as [169]
@@
$\mathcal{L}_{q}=-\frac{1}{2}\left(\nabla\phi\right)^{2}-V\left(\phi\right).$ (3)
Following the least action principle, we obtain the field equation in the form of [170]
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R+g_{\mu,\nu}\Lambda=-8\pi\bigg{(}\overline{T}_{\mu\nu}+T_{\mu\nu}\left(PFDM\right)-T_{\mu\nu}\left(q\right)\bigg{)}.$ (4)
@@
Since we consider dark matter as a type of perfect fluid, we can formulate the energy-momentum tensor in terms of energy density and pressure as
$T_{\mu}^{\nu}=\left[-\rho,p,p,p\right],$ (5)
In the simplest scenario, the ratio of energy density to pressure is assumed to have a constant value, $\frac{\rho}{p}=\delta-1$, [35]. The spherical symmetric solution of the field equations gives the Schwarzschild black hole solution in the form of [42, 170]
$ds^{2}=-f\left(r\right)dt^{2}+\frac{1}{f\left(r\right)}dr^{2}+r^{2}d\Omega^{2},$ (6)
with the lapse function, $f\left(r\right)$,
$f\left(r\right)=1-\frac{2M}{r}-\frac{\sigma}{r^{3\omega_{q}+1}}+\frac{\alpha}{r}\ln\frac{r}{\left|\alpha\right|}.$ (7)
Here, $\alpha$ is the PFDM intensity, and $\sigma$ and $\omega_{q}$ are the normalization factor and the quintessence state parameters, respectively.
@@
We notice that the quintessence state parameter, which varies between $-1/3<\omega_{q}<-1$, has a crucial impact on the lapse function. Therefore we have to discuss black hole properties with specific values of it. To be consistent throughout the rest of the manuscript, we choose the following four values: $\omega_{q}=-0.35$, $\omega_{q}=-0.55$, $\omega_{q}=-0.75$, and $\omega_{q}=-0.95$. At first in Fig. 1 we depict the lapse function with these parameters versus radius.
(a) $\omega_{q}=-0.35$.
(b) $\omega_{q}=-0.55$.
(c) $\omega_{q}=-0.75$.
@@
Figure 1: The impact of the dark matter and dark energy on the lapse function of Schwarzschild black hole for $\sigma=0.1$ and $M=1$.
We observe that the quintessence state parameter alters the concavity of the lapse function which is important in the context of the event horizon and naked singularity. For example, for $\omega_{q}=-0.35$ there is only one horizon. In this case, the dark matter parameter does not significantly influence the horizon radius. However, for smaller quintessence state parameters the concavity increases. For instance, for $\omega_{q}=-0.55$, we observe two horizons. Here, we note that for greater $\alpha$ values, the greater horizon radii take slightly greater values. For much smaller quintessence matter parameters, we see this impact more apparently. For example, for $\omega_{q}=-0.75$, the greater horizon radii occur relatively at smaller horizon values because of the larger concavity. All these cases are free of a naked singularity. However, for much smaller quintessence state parameters, for instance, for $\omega_{q}=-0.95$, the concavity increases so much that the lapse function may not reach zero for any radius. In this case, naked singularity can appear. On the other hand, with greater dark matter parameters an event horizon can occur and the naked singularity problem can be resolved. It is worth noting that these generic scenarios are plotted for a random mass value, $M=1$. For a different mass value, it is possible that there is no naked singularity.
@@
Fig. 2 shows that the mass function monotonously increases versus the event horizon. This finding is compatible with the presence of one event horizon. Here, we have to note that an increase in the dark matter parameter slightly affects the increase of the mass function. For smaller quintessence state parameters we observe that the monotonic increase of the mass function is no longer valid. In these cases, we observe that there is a lower limit as well as an upper limit on the event horizon. In Fig. 2, where $\omega_{q}=-0.55$, we hardly can detect these lower limit values since the upper limit values are much more greater than them. However, for $\omega_{q}=-0.75$ and $\omega_{q}=-0.95$, these lower bounds and intervals become apparent. In both cases, we observe a shift in the gap length as the impact of the dark matter parameter variation. To be more precise, we numerically calculate the lower and upper bound event horizon radii for these scenarios and present them in Table 1.

### v2

@@ Abstract
In this paper, we present an exact solution for a spherically symmetric Schwarzschild black hole surrounded by perfect fluid dark matter (PFDM) in the presence of a quintessence field. We investigate the impact of dark matter on the black hole’s thermodynamic and optical properties, as well as its quasinormal modes. Our analysis reveals a critical radius where the heat capacity becomes positive, indicating the thermodynamic stability of the black hole. Notably, this critical radius increases with the dark matter parameter $\alpha$. Additionally, we find that as the effects of dark matter increase, the black hole’s shadow radius decreases. Using the WKB approximation, we show that the quasinormal mode spectrum differs from that of a standard Schwarzschild black hole due to the influence of PFDM. Moreover, we demonstrate that as $\alpha$ increases, both the real part and the magnitude of the imaginary part of the quasinormal mode frequencies increase. This suggests that field perturbations decay more rapidly in the presence of PFDM compared to those in a Schwarzschild black hole.
@@ 1 Introduction
[…]
@@
Reliable astronomical observations put forward that our universe is expanding at an accelerating rate [51, 52, 53], and dark energy, the other component of the dark sector, is thought to be behind this fact. The cosmological constant, whose observational and theoretical values differ by $120$ orders of magnitude [54], cannot explain this expansion. Theoretical physicists proposed new models using dynamical scalar fields to describe dark energy, taking into account that the cosmological constant does not alter with time [55, 61, 56, 57, 58, 59, 60, 62]. Among different models, the quintessence matter model stands in the form where the equation of state is limited by the linear relationship between pressure, energy density and a state parameter [57]. Using the quintessence matter field, Kiselev solved Einstein’s equation and obtained a general solution form of the static spherically symmetric black hole [63]. Subsequently, in the presence of quintessence matter new solutions of other black holes were obtained, and the impacts of the quintessence field on their thermal quantities, shadows, and quasinormal modes were discussed widely in [64, 65, 66, 67, 68, 70, 71, 73, 72, 74, 75, 76, 78, 77, 80, 79, 83, 81, 82, 84, 85, 86, 87, 88, 90, 89, 91, 92, 69].
@@
Although black hole solutions have been sought for more than a century, the problem of their detection remained an open debate. Until the successful observational results of the Event Horizon Telescope were published [93, 94, 95, 96, 97], some authors suggested that not the black holes but their shadows could be detected [98, 100, 99]. According to this proposal, if a black hole is located between a light source and an observer, it should be expected that some of the light rays would be deflected by the intense gravitational field of the black hole and reach the observer. However, some photons would fall into the black hole instead of this deflection and create a dark region called a shadow. The apparent shape of the black hole is assumed to correspond to the boundary of the shadow. By studying black hole shadows, we can probe the effects of both dark matter and dark energy on the geometry of space-time, making them important tools for investigating alternative gravity theories and understanding the universe’s large-scale structure. With this motivation nowadays many papers examine the impact the space-time geometry, dark energy, and dark matter on shadow images in the literature [101, 102, 103, 109, 110, 125, 108, 113, 116, 121, 114, 112, 117, 118, 115, 122, 133, 128, 134, 119, 111, 104, 105, 106, 126, 127, 107, 129, 130, 140, 144, 131, 135, 141, 137, 120, 138, 139, 142, 132, 146, 143, 145, 123, 124, 136].
The detection of gravitational waves [147, 148] has provided significant insights into fundamental astrophysical phenomena [149]. Advanced observations of these waves have become essential tools for testing various theories that extend beyond general relativity, including both modified and alternative theories of gravity. Modified theories, such as $f(R)$ gravity [150, 151], Scalar-Tensor theory [152, 153], Einstein–Gauss-Bonnet theory [154], massive gravity theory [155] etc. stem from adjustments to the Einstein-Hilbert action while maintaining a geometrical structure consistent with general relativity. In contrast, alternative theories, including teleparallel gravity [156, 157, 158, 159], loop quantum gravity [160], string theory [161] introduce fundamentally different geometrical frameworks and offer unique predictions for gravitational waves, particularly under extreme conditions near black holes. Of particular interest, black holes in their final stages are theorized to emit gravitational waves as quasinormal radiation [162]. Mathematically, quasinormal modes are represented as complex numbers, with the real part corresponding to the oscillation frequency and the imaginary part representing the inverse damping time [163]. Their independence from initial perturbations allows them to be considered distinctive signatures of black holes [164]. A substantial body of literature has utilized both analytical and numerical methods to explore the quasinormal modes of black holes in various scenarios, thereby elucidating their implications for gravitational theories and advancing our understanding of black hole dynamics within the context of contemporary astrophysical observations [165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194].
Furthermore, the thermodynamics of black holes has become a fundamental area of study, particularly following the establishment of the connection between a black hole’s event horizon and its entropy [195, 196]. The presence of PFDM and quintessence fields modifies the energy-momentum tensor, altering the geometry of spacetime around black holes and impacting their thermodynamic properties, such as temperature, heat capacity, and introduce new critical phenomena. These modifications could lead to generalizations of classical thermodynamic results, contributing to our understanding of the information paradox and potentially offering new perspectives on holography and quantum gravity .
In conclusion, studying the quasinormal modes, shadows, and thermodynamics of black holes is crucial for investigating the influence of dark matter and dark energy, testing alternative gravity theories, and deepening our understanding of black holes’ fundamental properties and behavior under diverse cosmic conditions. Motivated by these consideration, this contribution aims to explore the thermodynamics, shadows, and quasinormal modes of a Schwarzschild black hole surrounded by PFDM, with the presence of the quintessence matter field. To achieve these goals, the paper is organized as follows: In Sec. 2, the mathematical framework of the scenario under consideration is briefly formulated. In Sec. 3, the thermodynamics of the system is examined. The anticipated shadow images and quasinormal modes of the black hole are then presented in Secs. 4 and 5, respectively, using two distinct approaches. Finally, the manuscript is concluded in Sec. 6. Note that throughout the manuscript, natural units are used, with $\hbar=k_{B}=1$.
@@ 2 Black Hole Solution
In the context of quintessence, the action associated with Einstein gravity in the presence of PFDM can be written as [35, 36, 42, 197]:
$S=\int d^{4}x\sqrt{-g}\left[{\frac{c^{4}}{16\pi G}R}-\mathcal{L}^{(PFDM)}+\mathcal{L}^{\left({quint}\right)}\right],$ (1)
@@
where $G$ is Newton’s constant, $c$ is the speed of light, $R$ is the scalar curvature, $\mathcal{L}^{(PFDM)}$ is the PFDM Lagrangian density, and $\mathcal{L}^{(quint)}$ is the Lagrangian density of the quintessence fields, which is expressed in terms of the quintessential scalar field $\phi$ and the potential $V\left(\phi\right)$ as follows [198]:
$\mathcal{L}^{\left({quint}\right)}=-\frac{1}{2}\left(\nabla\phi\right)^{2}-V\left(\phi\right).$ (2)
Applying the variational principle to Eq. (1) with respect to the contravariant metric tensor $g^{\mu\nu}$, we get
@@
$\frac{1}{\sqrt{-g}}\frac{\delta S}{\delta g^{\mu\nu}}=\frac{1}{\sqrt{-g}}\left[\frac{c^{4}}{4\pi G}\int d^{4}x\frac{\delta\left(\sqrt{-g}R\right)}{\delta g^{\mu\nu}}-\int d^{4}x\frac{\delta\left(\sqrt{-g}\mathcal{L}^{(PFDM)}\right)}{\delta g^{\mu\nu}}+\int d^{4}x\frac{\delta\left(\sqrt{-g}\mathcal{L}^{\left({quint}\right)}\right)}{\delta g^{\mu\nu}}\right]=0,$ (3)
which leads to the field equation
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\frac{8\pi G}{c^{4}}\Big{(}T_{\mu\nu}^{(PFDM)}-T_{\mu\nu}^{\left({quint}\right)}\Big{)},$ (4)
where the energy-momentum tensors are
$T_{\mu\nu}^{(PFDM)}$ $=$ $\frac{2\delta\left(\sqrt{-g}\mathcal{L}^{(PFDM)}\right)}{\delta g^{\mu\nu}},$ (5)
$T_{\mu\nu}^{\left({quint}\right)}$ $=$ $\frac{2\delta\left(\sqrt{-g}\mathcal{L}^{\left({quint}\right)}\right)}{\delta g^{\mu\nu}},$ (6)
@@
in units of $\frac{4\pi G}{c^{4}}=1$. Since we consider dark matter as a type of perfect fluid, we can represent it with an effective energy-momentum tensor. For a perfect fluid, the energy-momentum tensor is:
$T^{\mu\nu(PFDM)}=\left(\rho+p\right)u^{\mu}u^{\nu}+pg^{\mu\nu},$ (7)
where $\rho$ is the energy density, $p$ is the pressure, and $u^{\mu}$ is the four-velocity. In addition, for the simplest case, we assume [35],
$T_{\text{ \ }\theta}^{\theta(PFDM)}=T_{\text{ \ }\varphi}^{\varphi(PFDM)}=\left(1-\delta\right)T_{\text{ \ }t}^{t(PFDM)},$ (8)
where $\delta$ is a constant. To find a solution that satisfies Eq. (4), we assume a spherically symmetric line element [42, 199]:
$ds^{2}=-f\left(r\right)dt^{2}+\frac{1}{f\left(r\right)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right),$ (9)
or equivalently,
$ds^{2}=-\left(1-m\left(r\right)\right)dt^{2}+\frac{1}{1-m\left(r\right)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right).$ (10)
Replacing the metric ansatz in Eq. (4), the different components of Einstein equations take the form
$G_{0}^{0}=-\frac{m\left(r\right)}{r^{2}}-\frac{1}{r}\frac{dm\left(r\right)}{dr}+2T_{0}^{0\left({quint}\right)}=2T_{0}^{0\left({PFDM}\right)},$ (11)
$G_{2}^{2}=-\frac{1}{r}\frac{dm\left(r\right)}{dr}-\frac{1}{2}\frac{d^{2}m\left(r\right)}{dr^{2}}+2T_{2}^{2\left({quint}\right)}=2T_{2}^{2\left({PFDM}\right)},$ (12)
where
$T_{\mu}^{\nu\left({PFDM}\right)}=diag\left(-\rho,p_{r},p,p\right),$ (13)
and
$T_{0}^{0\left({quint}\right)}=T_{r}^{r\left({quint}\right)}=-\frac{3\sigma\omega_{q}}{2r^{3\left(\omega_{q}+1\right)}};T_{2}^{2\left({quint}\right)}=-\frac{\left(3\omega_{q}+1\right)}{2}T_{0}^{0\left({quint}\right)}.$ (14)
Furthermore, in the simplest case, we consider the equation of state for the PFDM as
$p=\left(\delta-1\right)\rho,$ (15)
and taking the ratio of the above Eqs. (11) and (12), we get
$\frac{1}{2}\frac{d^{2}m\left(r\right)}{dr^{2}}+\frac{\delta}{r}\frac{dm\left(r\right)}{dr}+\left(\delta-1\right)\frac{m\left(r\right)}{r^{2}}-\frac{\left(3\omega_{q}+1\right)}{2}\frac{3\sigma\omega_{q}}{r^{3\left(\omega_{q}+1\right)}}+\left(\delta-1\right)\frac{3\sigma\omega_{q}}{r^{3\left(\omega_{q}+1\right)}}=0.$ (16)
The solution of this equation for $\delta=3/2$ is
$m\left(r\right)=\frac{r_{s}}{r}-\frac{\alpha}{r}\ln\frac{r}{\left|\alpha\right|}+\frac{\sigma}{r^{1+3\omega_{q}}},$ (17)
where $r_{s}$ and $\alpha$ are the integration constants. To find $r_{s}$, we consider the case where both $\alpha=\sigma=0$. Under these conditions, we find that $r_{s}=2M$. This then determines the form of the lapse function,
$f\left(r\right)=1-\frac{2M}{r}+\frac{\alpha}{r}\ln\frac{r}{\left|\alpha\right|}-\frac{\sigma}{r^{1+3\omega_{q}}}.$ (18)
@@
We notice that the quintessence state parameter, which varies between $-1/3<\omega_{q}<-1$, has a crucial impact on the metric function. Therefore we have to discuss black hole properties with specific values of it. To be consistent throughout the rest of the manuscript, we choose the following four values: $\omega_{q}=-0.35$, $\omega_{q}=-0.55$, $\omega_{q}=-0.75$, and $\omega_{q}=-0.95$. At first in Fig. 1 we depict the metric function with these parameters versus radius.
@@
Figure 1: The impact of the dark matter and dark energy on the metric function of Schwarzschild black hole for $\sigma=0.1$ and $M=1$.
We observe that the quintessence state parameter alters the concavity of the metric function which is important in the context of the event horizon and naked singularity. For example, for $\omega_{q}=-0.35$ there is only one horizon. In this case, the dark matter parameter does not significantly influence the horizon radius. However, for smaller quintessence state parameters the concavity increases. For instance, for $\omega_{q}=-0.55$, we observe that as PFDM influence increases, the metric function $f(r)$ peaks higher and decreases more slowly with distance. All curves indicate two horizons where $f(r)=0$. Stronger PFDM effects (higher $\alpha$) appear to expand the region between these horizons and extend the black hole’s gravitational influence farther into space. The graph suggests dark matter significantly alters the black hole’s spacetime structure, potentially affecting its observable properties. For much smaller quintessence matter parameters, we see that the impact of PFDM more apparently. For example, for $\omega_{q}=-0.75$, the greater horizon radii occur relatively at smaller horizon values because of the larger concavity. All these cases are free of a naked singularity. However, for much smaller quintessence state parameters, for instance, for $\omega_{q}=-0.95$, the concavity increases so much that the metric function may not reach zero for any radius. In this case, naked singularity can appear. On the other hand, with greater dark matter parameters an event horizon can occur and the naked singularity problem can be resolved. It is worth noting that these generic scenarios are plotted for a random mass value, $M=1$. For a different mass value, it is possible that there is no naked singularity.
@@
Fig. 2, shows that the black hole mass grows as $\alpha$ increases, for the same horizon size, indicating that dark matter may enhance the black hole’s mass. The nearly linear behavior of the curves suggests a proportional relationship between mass and horizon size, with a slight deviation at smaller radii that could be due to additional dark matter effects. In these cases, we observe that there is a lower limit. In Figs. 2, 2, and 2, we see that all curves exhibit a similar parabolic shape, starting near the origin. As $\alpha$ increases, the curves reach higher peak masses, suggesting that a higher dark matter content allows for more massive black holes. After reaching a maximum, the mass begins to decrease until it approaches zero. When the mass approaches its maximum, the black hole may undergo a change in stability, This extremum of the Hawking mass can obtained from $\frac{\partial}{\partial r_{H}}M=0.$. In all these cases, we observe that there is a lower limit as well as an upper limit on the event horizon. To be more precise, we numerically calculate the lower and upper bound event horizon radii for these scenarios and present them in Table 1.